OpenStudy (anonymous):

Can you show me how to do solving two step questions and a check of the problem. 3x + 1 = 10 or any problem so I can learn how to do it :D Thanks!

OpenStudy (anonymous):

3x+1 = 10 3x=9 then you divide 9 by 3 so you can keep X on one side So the answer is 3 CHECK---------\ 3x+1=10 plug in 3 for x 3*3 +1 = 10 9+1=10 10=10

OpenStudy (anonymous):

Thank you :D Giving you award right now. <3

OpenStudy (anonymous):

You basically want to isolate X. Then once you find what X is = to, plug the value of X that you just found into the original equation. Later you will just ignore the check part but if you are unsure, you can always check it by plugging it in.